OpenStudy (anonymous):

Divide the polynomials by the monomials. (21y4 + 18y3) /3y

OpenStudy (anonymous):

Consider, \[\frac{ 21y^4+18y^3 }{ 3y } \implies \frac{ 21y^4 }{ 3y }+\frac{ 18y^3 }{ 3y }\]